The directions to the question: find the equation of the line parallel to the
y=4x-2 that passes through the point (0,4). I don't know how to find the answer to this problem. Please help and thank you for your time.

Answer by oscargut(2103)

let the equation to find y=ax+b
is parallel to the y=4x-2 so a=4
so y=4x+b but passes through the point (0,4) then 4=4(0)+b then b=4
then the solution is y=4x+4